# Offline Goal-Conditioned Reinforcement Learning with Projective Quasimetric Planning

### v2

@@ 1 Introduction
Humans improve knowledge and master complex skills by iteratively exploring and refining internal notions of proximity ("How far am I from success ?") and progress ("Is this action helpful ?") (practice, ; internal, ). Reinforcement learning (RL) (rl, ) offers the computational analogue, combining trial-and-error with value function approximation to learn algorithms and agents to solve sequential decision problems, extended to high-dimensional perception and control with Deep Reinforcement Learning (DRL) (deeprl, ).
@@
To build AI agents in game productions, live exploration is costly ; however, vast logs of diverse human play exist. Offline RL addresses data-efficiency concerns by learning from a fixed set of trajectories collected a priori (offlinerl, ). Without the ability to query new actions, the agents must optimize their policy, eventually beyond the dataset behavior distribution iql ; awr . Nevertheless, accurately estimating long-horizon value functions is notably difficult as bootstrapping errors may occur (h_iql, ).
@@
By conditioning an agent’s policy on a target outcome, Goal-Conditioned RL (GCRL) repurposes the sequential decision problem as a directed reachability task (gcrlsurvey, ; gcrl1993, ). Although techniques like Hindsight Experience Replay (HER) her mitigate sparse rewards and Hierarchical Policies (HP) h_iql ; h_gcoff ease path planning, most methods fall short when tackling complex long-horizon navigation problems, which is our main focus in our research. Recent benchmarks like D4RL (d4rl, ) and OGBench (ogbench, ) highlight these difficulties within diverse environments and multiple dataset distributions.
@@
To tackle these failures, a relevant remedy to consider is to simplify the planning problem by learning a latent representation that captures reachability. Contrastive and Self-Supervised (curl, ; ssgcrl, ) methods cluster similar states but lack calibrated distances ; Bisimulation metrics and MDP abstractions(bisim, ; deepmdp, ) preserve dynamics but involve variational objectives and offer no out-of-distribution guard ;Current landmark-based mapping methods pre-compute distances to fixed keypoints but may suffer coverage gaps (mapping, ) ; Recently, HILP (hilp, ) took an interesting geometric turn by embedding states into a Hilbert space where straight-line length, or Euclidean norm, approximates time-to-reach. However, at test time, it must scan the entire offline dataset for nearest neighbors, and its enforced symmetry overlooks the inherently directional nature of goal reachability.
@@
Directionality matters : descending a cliff is simple ; climbing back is not. Quasimetric RL (QRL) (iqe, ; qrl, ) embraces this asymmetry, learning a quasimetric (non symmetric distance) upper-bounding transitions to an unit cost while pushing apart non-successor states, and consequently approximates the minimum reach time between any states. Nonetheless, the policy learning is brittle as it only relies on model-based rollouts, which fail in long-horizon settings as inaccuracies accumulate.
@@
We use a goal-conditioned policy trained with Advantage Weighted Regression (AWR) awr to generate actions toward nearby latent targets, a simpler learning task than direct long-range navigation. For distant goals, we employ efficient Floyd-Warshall lookups across keypoints. This approach avoids distance symmetry constraints of HILP and scales memory independently of dataset size.
@@ 2 Related Work
As the name implies, Offline GCRL lies at the intersection of Offline RL (batchrl, ; offlinerl, ) and GCRL gcrl1993 ; gcrlsurvey , and aims to address two main challenges : trajectory stitching and credit assignment (d4rl, ; ogbench, ). Prior works tackled those issues by leveraging relabeling methods. For instance, gcsl ; wgcsl apply Hindsight Experiment Replay (HER) her with Imitation Learning. Leveraging this basis, actionablemodels ; iql additionally performs value learning and policy extraction through Offline Q-Learning. Nevertheless, both kinds of methods strugglein long-range scenarios, notably when the value estimates suffer from a low signal-to-noise ratio.
@@
Decomposing the decision-making by leveraging a hierarchical structure is a long-standing idea (schmidhuber1991learning, ; sutton1999between, ). Recent work showed its interest in solving long-horizon tasks in online settings (feudalhnets, ; ris, ; hdrl, ; lmlhh, ; dataeff, ; nearopt, ; gacshrl, ) and in offline ones (h_iql, ; qphil, ). While many Hierarchical GCRL (HGCRL) methods relate to the offline skill extraction literature opal ; tap ; ddco ; acrllsp ; tacorl ; sbmbrl ; por ; h_iql ; hilp ; hilow , which aims to decompose the goal-reaching task into a succession of sub-tasks (as trajectory segments, directions, or precise subgoals), our method relates more to the graph planning strategies. Indeed, learning graphs for planning has been broadly used in the online setting to alleviate exploration problems sorb ; sfllggcrl ; msplugr ; lgsghrl ; igbpgcp ; pgcp ; sptmn ; wmglllp ; tpe ; dhrl ; cqm , unlike ProQ, which targets the offline setting. Recent work started to use graph learning as a way to ease long-range planning in the offline settings (vmg, ; qphil, ), however without relying on a quasimetric distance, which allow in our framework the learning of a uniformly spanning set of in-distribution keypoints.
@@
Leveraging temporally meaningful representations for planning in a complex state space has been a growing research topic in GCRL. While several prior works relied on contrastive losses to encode temporal similarity within the representations (crl, ; qphil, ; plan2vec, ; minedojo, ; vmg, ; r3m, ), they did so without a real metric calibration, as ProQ does. Instead of a contrastive approach, our method leverages a distance function within a latent state space to allow for metric-grounded planning. Likewise, Foundation Policies with Hilbert Representations (HILP) hilp builds a Hilbert representation space to encode temporal similarity of states within its induced metric, but without taking into account asymmetry like our framework does. Additionally, Quasimetric Reinforcement Learning (QRL) qrl , which is closely related to our proposed method, leverages Interval Quasimetric Embeddings (IQE) iqe to learn an asymmetric distance-value function. However, while QRL uses it to extract a policy through a valued model-based approach, we leverage it to guide the learning of a uniformly distributed set of keypoints to perform long-range planning.
@@ 3 Preliminaries
Unlike traditional metrics, which are symmetric, an optimal value function in GCRL often exhibits asymmetry as an optimal path from a state A to a state B is may not be the same to go from B to A. This property aligns with the concept of a Quasimetric onquasi , a function $d:\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}^{+}$ satisfying three rules : Non-negativity : $d(x,y)\geq 0$ ; Identity of indiscernibles : $d(x,x)=0$ ; Triangle inequality : $d(x,z)\leq d(x,y)+d(y,z)$ ; but not necessarily symmetry : $d(x,y)\neq d(y,x)$.
To model such a distance, we consider : Interval Quasimetric Embeddings (IQE) (iqe, ), and its continuation, Quasimetric Reinforcement Learning (QRL) (qrl, ). IQE represents the quasimetric as the aggregated length of intervals in a latent space and is a universal approximation for distances.
@@
$\forall\leavevmode\nobreak\ i=1,\ldots,N\ \ \ \ \ \ \ \ \ \ d_{\theta^{\prime}}^{(i)}(x,y)=\Bigg{|}\bigcup_{j=1}^{M}\Big{[}f_{\theta^{\prime}}(x)_{ij},\max\big{(}f_{\theta^{\prime}}(x)_{ij},f_{\theta^{\prime}}(y)_{ij}\big{)}\Big{]}\Bigg{|}$ (1)
IQE components are positive homogeneous and can be arbitrarily scaled, and thus do not require special reparametrization in combining them. Using the maxmean reduction from prior work (pitis, ), we obtain IQE-maxmean with a single extra trainable parameter $\alpha\in[0,1]$, and considering $\theta=[\theta^{\prime},\alpha]$ :
@@
IQE Universal Approximation General Case (See (iqe, ) for the demonstration) : Consider any quasimetric space $(\mathcal{X},d)$ where $\mathcal{X}$ is compact and $d$ is continuous. $\forall\leavevmode\nobreak\ \epsilon>0$,- with sufficiently large $N$, there exists a continuous encoder $f_{\theta^{\prime}}:\mathcal{X}\rightarrow\mathbb{R}^{N\times M}$ and $\alpha\in[0,1]$ such that :
$\forall x,y\in\mathcal{X},\leavevmode\nobreak\ \big{|}d_{\theta}(x,y)-d(x,y)\big{|}\leq\epsilon\leavevmode\nobreak\ .$
@@
QRL Function Approximation [General,Formal] (See qrl for the demonstration) :Consider a compact state space $\mathcal{S}$ and an optimal value function $V^{*}$. We suppose that the state space is the goal space ($\mathcal{S}=\mathcal{G}$). If $\{d_{\theta}\}_{\theta\in\Theta}$ are universal approximators of quasimetrics over $\mathcal{S}$ (e.g. IQE) :
$\forall\epsilon>0,\leavevmode\nobreak\ \exists\theta^{*}\in\Theta,\forall s,g\in\mathcal{S},\leavevmode\nobreak\ \mathbb{P}\Big{[}\big{|}d_{\theta^{*}}(s,g)+(1+\epsilon)V^{*}(s,g)\big{|}\in\big{[}-\sqrt{\epsilon},0\big{]}\Big{]}=1-\mathcal{O}\big{(}-\sqrt{\epsilon}\cdot\mathbb{E}[V^{*}]\big{)}$
i.e. $d_{\theta^{*}}$ recovers $-V^{*}$ up to a known scale, with probability $1-\mathcal{O}\big{(}-\sqrt{\epsilon}\cdot\mathbb{E}[V^{*}]\big{)}$
@@
$V(x_{1},x_{2})=\frac{q_{1}q_{2}}{4\pi\epsilon_{0}}\cdot\frac{1}{\lVert x_{1}-x_{2}\rVert_{2}}\leavevmode\nobreak\ ,$
@@
$E(x_{1},\ldots,x_{K})=\sum_{1\leq i<j\leq K}\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\leavevmode\nobreak\ .$
Minimizing this energy $E(x_{1},\ldots,x_{K})$ on a compact domain (e.g. the unit sphere $\mathbb{S}^{2}$) yields the classical Thomson Problem (thompsonsphere, ; thompsonatomic, ) of finding well-separated, nearly uniform point configurations.
@@ 4.1 ProQ Learning Algorithm : Overview
1:initial state $s_{0}\in\mathcal{S}$ , goal $g\in\mathcal{G}$ , latent keypoints $\{z_{k}\}_{k=1}^{K}$ , maximum horizon $T$
2: Init: Build  $G=\big{\{}z_{k}\big{\}}_{k=1}^{K+1}$ with $z_{k+1}=\phi(g)$
3: Init: Compute $D^{*}=Distances^{*}\big{(}\{z_{k}\}_{k=1}^{K+1}\big{)}\leftarrow$ Floyd–Warshall$\big{(}G\big{)}$ ;
@@
9:     $s\!\leftarrow$ env.step$(a)$ ;
10:     $t\leftarrow t+1$
@@ 4.2 Learning a Robust Latent Space
To obtain a structured latent space that captures meaningful differences between states, we encode each observation $s$ into a latent vector $z=\phi(s)$ using a neural encoder $\phi$ trained with VICReg (vicreg, ). VICReg balances three terms : (1) an invariance loss aligning paired embeddings ; (2) a variance regularization ensuring non-collapsed features; and (3) a decorrelation loss to prevent redundancy. We consider the variance and covariance regularizers :
$\small\mathcal{L}_{\text{var}}(X)=\frac{1}{D}\sum_{i=1}^{D}\max\Big{(}0,\gamma-\sqrt{\text{Var}(x^{i})}\Big{)}$ (4)
$\small\mathcal{L}_{\text{cov}}=\frac{1}{D}\sum_{i\neq j}C(X)_{ij}^{2}$ (5)
@@
  • Extrapolated states : $\tilde{z}^{e}=\phi\bigl{(}(1+\beta)\cdot s_{1}-\beta\cdot s_{2}\bigr{)}$, $\beta\sim\mathcal{U}([0,1])$
Using these generated states, we learn $\psi$ with a binary cross-entropy with a Lagrange multiplier to enforce high confidence on positives while pushing negatives in-distribution probability towards 00 :
@@
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\Big{(}(1-\delta)-\psi(z)\Big{)}^{2}$, with $10<\delta<$ 1 the compliance.
@@
In/Out-of-Distribution Generalization Guarantee (See Appendix D.1.1 for proof) : Let $\mathcal{D}$ be an $\epsilon$-dense dataset of points within a compact state space $\mathcal{S}\subset\mathbb{R}^{n}$. We consider $\widetilde{\mathcal{D}}$, a set of generated points from $\mathcal{D}$, which is $\epsilon$-dense in the shell $\bigl{\{}\ \tilde{s}\in\mathbb{R}^{n}\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \text{dist}(\tilde{s},\mathcal{D})\leq Diam(\mathcal{D})\leavevmode\nobreak\ \bigr{\}}$.After convergence of the latent space component $\phi$, $\psi$, and $d$, considering the compliance $10<\delta<$ 1, there exist constants $10<C,0<\eta<\mathrm{Diam}(\mathcal{D}),0<\bar{\delta}<$ 1 such that:

### v3

@@ 1 Introduction
Humans improve knowledge and master complex skills by iteratively exploring and refining internal notions of proximity ("How far am I from success ?") and progress ("Is this action helpful ?") (Ericsson et al., 1993; Kawato, 1999). Reinforcement learning (RL) (Sutton and Barto, 2018) offers the computational analogue, combining trial-and-error with value function approximation to learn algorithms and agents to solve sequential decision problems, extended to high-dimensional perception and control with Deep Reinforcement Learning (DRL) (Arulkumaran et al., 2017).
@@
To build AI agents in game productions, live exploration is costly ; however, vast logs of diverse human play exist. Offline RL addresses data-efficiency concerns by learning from a fixed set of trajectories collected a priori (Levine et al., 2020). Without the ability to query new actions, the agents must optimize their policy, eventually beyond the dataset behavior distribution Kostrikov et al. (2021); Peng et al. (2019). Nevertheless, accurately estimating long-horizon value functions is notably difficult as bootstrapping errors may occur (Park et al., 2023).
@@
By conditioning an agent’s policy on a target outcome, Goal-Conditioned RL (GCRL) repurposes the sequential decision problem as a directed reachability task (Liu et al., 2022; Kaelbling, 1993). Although techniques like Hindsight Experience Replay (HER) Andrychowicz et al. (2017) mitigate sparse rewards and Hierarchical Policies (HP) Park et al. (2023); Li et al. (2022) ease path planning, most methods fall short when tackling complex long-horizon navigation problems, which is our main focus in our research. Recent benchmarks like D4RL (Fu et al., 2020) and OGBench (Park et al., 2024a) highlight these difficulties within diverse environments and multiple dataset distributions.
@@
To tackle these failures, a relevant remedy to consider is to simplify the planning problem by learning a latent representation that captures reachability. Contrastive and Self-Supervised (Laskin et al., ; Venkattaramanujam et al., 2019) methods cluster similar states but lack calibrated distances ; Bisimulation metrics and MDP abstractions(Shimizu and Tomizuka, 2024; Gelada et al., ) preserve dynamics but involve variational objectives and offer no out-of-distribution guard ;Current landmark-based mapping methods pre-compute distances to fixed keypoints but may suffer coverage gaps (Huang et al., 2019a) ; Recently, HILP (Park et al., 2024b) took an interesting geometric turn by embedding states into a Hilbert space where straight-line length, or Euclidean norm, approximates time-to-reach. However, at test time, it must scan the entire offline dataset for nearest neighbors, and its enforced symmetry overlooks the inherently directional nature of goal reachability.
@@
Directionality matters : descending a cliff is simple ; climbing back is not. Quasimetric RL (QRL) (Wang and Isola, 2022; Wang et al., 2023) embraces this asymmetry, learning a quasimetric (non symmetric distance) upper-bounding transitions to an unit cost while pushing apart non-successor states, and consequently approximates the minimum reach time between any states. Nonetheless, the policy learning is brittle as it only relies on model-based rollouts, which fail in long-horizon settings as inaccuracies accumulate.
@@
We use a goal-conditioned policy trained with Advantage Weighted Regression (AWR) Peng et al. (2019) to generate actions toward nearby latent targets, a simpler learning task than direct long-range navigation. For distant goals, we employ efficient Floyd-Warshall lookups across keypoints. This approach avoids distance symmetry constraints of HILP and scales memory independently of dataset size.
@@ 2 Related Work
As the name implies, Offline GCRL lies at the intersection of Offline RL (Lange et al., 2012; Levine et al., 2020) and GCRL Kaelbling (1993); Liu et al. (2022), and aims to address two main challenges : trajectory stitching and credit assignment (Fu et al., 2020; Park et al., 2024a). Prior works tackled those issues by leveraging relabeling methods. For instance, Ghosh et al. (2019); Yang et al. (2022) apply Hindsight Experiment Replay (HER) Andrychowicz et al. (2017) with Imitation Learning. Leveraging this basis, Chebotar et al. (2021); Kostrikov et al. (2021) additionally performs value learning and policy extraction through Offline Q-Learning. Nevertheless, both kinds of methods strugglein long-range scenarios, notably when the value estimates suffer from a low signal-to-noise ratio.
@@
Decomposing the decision-making by leveraging a hierarchical structure is a long-standing idea (Schmidhuber, 1991; Sutton et al., 1999). Recent work showed its interest in solving long-horizon tasks in online settings (Vezhnevets et al., 2017; Chane-Sane et al., 2021; Kulkarni et al., 2016; Levy et al., 2019; Nachum et al., 2018, 2019; Zhang et al., 2021b) and in offline ones (Park et al., 2023; Canesse et al., 2024). While many Hierarchical GCRL (HGCRL) methods relate to the offline skill extraction literature Ajay et al. (2021); Jiang et al. (2023); Krishnan et al. (2017); Pertsch et al. (2020); Rosete-Beas et al. (2022); Shi et al. (2022); Xu et al. (2023); Park et al. (2023, 2024b); Kobanda et al. (2024), which aims to decompose the goal-reaching task into a succession of sub-tasks (as trajectory segments, directions, or precise subgoals), our method relates more to the graph planning strategies. Indeed, learning graphs for planning has been broadly used in the online setting to alleviate exploration problems Eysenbach et al. (2019); Hoang et al. (2021); Huang et al. (2019b); Kim et al. (2021, 2023); Nasiriany et al. (2019); Savinov et al. (2018); Zhang et al. (2021a); Hong et al. (2023); Lee et al. (2022, 2023), unlike ProQ, which targets the offline setting. Recent work started to use graph learning as a way to ease long-range planning in the offline settings (Zhu et al., 2023; Canesse et al., 2024), however without relying on a quasimetric distance, which allow in our framework the learning of a uniformly spanning set of in-distribution keypoints.
@@
Leveraging temporally meaningful representations for planning in a complex state space has been a growing research topic in GCRL. While several prior works relied on contrastive losses to encode temporal similarity within the representations (Eysenbach et al., 2023; Canesse et al., 2024; Yang et al., 2020; Fan et al., 2022; Zhu et al., 2023; Nair et al., 2022), they did so without a real metric calibration, as ProQ does. Instead of a contrastive approach, our method leverages a distance function within a latent state space to allow for metric-grounded planning. Likewise, Foundation Policies with Hilbert Representations (HILP) Park et al. (2024b) builds a Hilbert representation space to encode temporal similarity of states within its induced metric, but without taking into account asymmetry like our framework does. Additionally, Quasimetric Reinforcement Learning (QRL) Wang et al. (2023), which is closely related to our proposed method, leverages Interval Quasimetric Embeddings (IQE) Wang and Isola (2022) to learn an asymmetric distance-value function. However, while QRL uses it to extract a policy through a valued model-based approach, we leverage it to guide the learning of a uniformly distributed set of keypoints to perform long-range planning.
@@ 3 Preliminaries
Unlike traditional metrics, which are symmetric, an optimal value function in GCRL often exhibits asymmetry as an optimal path from a state A to a state B is may not be the same to go from B to A. This property aligns with the concept of a Quasimetric Wilson (1931), a function $d:\mathcal{G}\times\mathcal{G}\rightarrow\mathbb{R}^{+}$ satisfying three rules : Non-negativity : $d(x,y)\geq 0$ ; Identity of indiscernibles : $d(x,x)=0$ ; Triangle inequality : $d(x,z)\leq d(x,y)+d(y,z)$ ; but not necessarily symmetry : $d(x,y)\neq d(y,x)$.
To model such a distance, we consider : Interval Quasimetric Embeddings (IQE) (Wang and Isola, 2022), and its continuation, Quasimetric Reinforcement Learning (QRL) (Wang et al., 2023). IQE represents the quasimetric as the aggregated length of intervals in a latent space and is a universal approximation for distances.
@@
$\forall\penalty 10000\ i=1,\ldots,N\ \ \ \ \ \ \ \ \ \ d_{\theta^{\prime}}^{(i)}(x,y)=\Bigg|\bigcup_{j=1}^{M}\Big[f_{\theta^{\prime}}(x)_{ij},\max\big(f_{\theta^{\prime}}(x)_{ij},f_{\theta^{\prime}}(y)_{ij}\big)\Big]\Bigg|$ (1)
IQE components are positive homogeneous and can be arbitrarily scaled, and thus do not require special reparametrization in combining them. Using the maxmean reduction from prior work (Pitis et al., 2020), we obtain IQE-maxmean with a single extra trainable parameter $\alpha\in[0,1]$, and considering $\theta=[\theta^{\prime},\alpha]$ :
@@
IQE Universal Approximation General Case (See (Wang and Isola, 2022) for the demonstration) : Consider any quasimetric space $(\mathcal{X},d)$ where $\mathcal{X}$ is compact and $d$ is continuous. $\forall\penalty 10000\ \epsilon>0$,- with sufficiently large $N$, there exists a continuous encoder $f_{\theta^{\prime}}:\mathcal{X}\rightarrow\mathbb{R}^{N\times M}$ and $\alpha\in[0,1]$ such that :
$\forall x,y\in\mathcal{X},\penalty 10000\ \big|d_{\theta}(x,y)-d(x,y)\big|\leq\epsilon\penalty 10000\ .$
@@
QRL Function Approximation [General,Formal] (See Wang et al. (2023) for the demonstration) :Consider a compact state space $\mathcal{S}$ and an optimal value function $V^{*}$. We suppose that the state space is the goal space ($\mathcal{S}=\mathcal{G}$). If $\{d_{\theta}\}_{\theta\in\Theta}$ are universal approximators of quasimetrics over $\mathcal{S}$ (e.g. IQE) :
$\forall\epsilon>0,\penalty 10000\ \exists\theta^{*}\in\Theta,\forall s,g\in\mathcal{S},\penalty 10000\ \mathbb{P}\Big[\big|d_{\theta^{*}}(s,g)+(1+\epsilon)V^{*}(s,g)\big|\in\big[-\sqrt{\epsilon},0\big]\Big]=1-\mathcal{O}\big(-\sqrt{\epsilon}\cdot\mathbb{E}[V^{*}]\big)$
i.e. $d_{\theta^{*}}$ recovers $V^{*}$ up to a known scale, with probability $1-\mathcal{O}\big(-\sqrt{\epsilon}\cdot\mathbb{E}[V^{*}]\big)$ .
@@
$V(x_{1},x_{2})=\frac{q_{1}q_{2}}{4\pi\epsilon_{0}}\cdot\frac{1}{\lVert x_{1}-x_{2}\rVert_{2}}\penalty 10000\ ,$
@@
$E(x_{1},\ldots,x_{K})=\sum_{1\leq i<j\leq K}\frac{1}{\lVert x_{i}-x_{j}\rVert_{2}}\penalty 10000\ .$
Minimizing this energy $E(x_{1},\ldots,x_{K})$ on a compact domain (e.g. the unit sphere $\mathbb{S}^{2}$) yields the classical Thomson Problem (Brown, 2020; LaFave Jr, 2013) of finding well-separated, nearly uniform point configurations.
@@ 4.1 ProQ Learning Algorithm : Overview
1:initial state $s_{0}\in\mathcal{S}$ , goal $g\in\mathcal{G}$ , latent keypoints $\{z_{k}\}_{k=1}^{K}$ , maximum horizon $T$ .
2:Init: Build $G=\big\{z_{k}\big\}_{k=1}^{K+1}$ with $z_{k+1}=\phi(g)$ .
[…]
@@
9:  $s\!\leftarrow$ env.step(a)(a) ;
10:  $t\leftarrow t+1$ .
@@ 4.2 Learning a Robust Latent Space
To obtain a structured latent space that captures meaningful differences between states, we encode each observation $s$ into a latent vector $z=\phi(s)$ using a neural encoder $\phi$ trained with VICReg (Bardes et al., 2021). VICReg balances three terms : (1) an invariance loss aligning paired embeddings ; (2) a variance regularization ensuring non-collapsed features; and (3) a decorrelation loss to prevent redundancy. We consider the variance and covariance regularizers :
$\mathcal{L}_{\text{var}}(X)=\frac{1}{D}\sum_{i=1}^{D}\max\Big(0,\gamma-\sqrt{\text{Var}(x^{i})}\Big)$ (4)
$\mathcal{L}_{\text{cov}}=\frac{1}{D}\sum_{i\neq j}C(X)_{ij}^{2}$ (5)
@@
  • Extrapolated states : $\tilde{z}^{e}=\phi\bigl((1+\beta)\cdot s_{1}-\beta\cdot s_{2}\bigr)$, $\beta\sim\mathcal{U}([0,1])$ .
Using these generated states, we learn $\psi$ with a binary cross-entropy with a Lagrange multiplier to enforce high confidence on positives while pushing negatives in-distribution probability towards 0 :
@@
with $\lambda_{\text{ood}}$ updated until satisfaction of a penalty $\Big((1-\delta)-\psi(z)\Big)^{2}$, with $0<\delta<1$ the compliance.
@@
In/Out-of-Distribution Generalization Guarantee (See Appendix D.1.1 for proof) : Let $\mathcal{D}$ be an $\epsilon$-dense dataset of points within a compact state space $\mathcal{S}\subset\mathbb{R}^{n}$. We consider $\widetilde{\mathcal{D}}$, a set of generated points from $\mathcal{D}$, which is $\epsilon$-dense in the shell $\bigl\{\ \tilde{s}\in\mathbb{R}^{n}\penalty 10000\ \big|\penalty 10000\ \text{dist}(\tilde{s},\mathcal{D})\leq Diam(\mathcal{D})\penalty 10000\ \bigr\}$.After convergence of the latent space component $\phi$, $\psi$, and $d$, considering the compliance $0<\delta<1$, there exist constants $0<C,0<\eta<\mathrm{Diam}(\mathcal{D}),0<\bar{\delta}<1$ such that: